• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name pking

    Log entry time 23:59:34 on April16,2001

    Entry number 60675

    keyword=Swing shift summary.

    The accelerator has been in the process of recovering beam and setting up the new
    energy. There have been problems with the magnet power supplies in Arc 1, and
    the hall C extraction magnet. There has been no beam through the swing shift.

    MCC does not expect to get beam into the hall until late in owl shift, or perhaps
    even into day shift. They will update us as their status changes.

    Currently, we have the 9-foil carbon optics target in position. Once beam is on
    target, Pete should be called.